# Convergent Privacy Framework for Multi-layer GNNs through Contractive Message Passing

### v2

@@ I Introduction
Graph neural networks (GNNs) [iclr/KipfW17, nips/GuC0SG20, iclr/VelickovicCCRLB18], designed for operating over structural data, have achieved success in various domains, including social networks [kdd/PerozziAS14, www/ZhangGPH22] and recommendation systems [www/ChenBSXZHHWH24, sigir/WuWF0CLX21, csur/WuSZXC23]. At their core, many GNN architectures are built upon the message-passing paradigm, where node representations are iteratively updated by aggregating information from their neighbors. However, graph structures often encode sensitive information about relationships and attributes. As a result, GNNs are vulnerable to privacy attacks, including membership inference [sp/Wulinkteller22, ccs/0001MMBS22, corr/abs-2102-05429] and attribute inference [uss/000100S022, tkde/SunDYZWYHL23]. These vulnerabilities highlight the urgent need for robust privacy protection mechanisms in graph learning.
@@
Differential privacy (DP) [csf/mironov2017renyi, wasserman2010statistical, journal/dong2022gaussian] has emerged as a foundational framework to provide formal guarantees against data leakage over the past two decades, with widespread applications in machine learning [nips/TangPSM23, sp/Yu0PGT19, nips/YangZZ0PL023], synthetic data generation [imc/SunCGWL24, iclr/LiWC23, jpc/McKennaMS21, pvldb/PatwaSGMR23], and beyond. In the context of GNNs, recent works [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23] have advanced privacy protection through edge-level DP (EDP) [sp/0011L0022] and node-level DP (NDP) [pakdd/IftikharW21, sigmod/DayLL16] guarantees. The primary approach employs perturbed message passing, which injects calibrated Gaussian or Laplace noise into aggregation layers to protect the edge or node memberships in the training graph. While these approaches have provided formal privacy guarantees, they share a critical limitation: the privacy loss grows linearly with the number of layers $K$ or graph hops. In other words, each additional aggregation layer compounds the privacy cost, ultimately requiring large amounts of noise to maintain a reasonable level of privacy protection. This, in turn, severely degrades model utility.
@@
Recent advances have shown that multi-layer GNNs, especially deeper GCNs, are essential in capturing complex relationships [feng2022powerful] and analyzing graphs with long-range interactions [nips/DwivediRGPWLB22, corr/abs-2406-03386, nips/WuJWMGS21, tmlr/TonshoffRRG24, iclr/ThurlemannR23]. In fact, informative long-range interactions exist in a lot of real-world datasets. For instance, in large biological networks, long-range dependencies influence protein functions, requiring more than 10 hops of message passing [icml/Li0GK21]. In social networks, privacy-sensitive relationships propagate through multi-hop neighborhoods [nguyen2017probability, liu2014assessment]. As reported in [icml/Li0GK21], increasing the network depth leads to a substantial improvement in accuracy from $72.5\%$ to $88.2\%$. However, the aforementioned linear dependence on $K$ is particularly challenging for multi-layer GNNs as larger $K$ leads to larger privacy parameter $\epsilon$, a.k.a weak privacy guarantee.
@@
Interestingly, empirical studies [sp/Wulinkteller22] have shown that membership inference attacks are not particularly more successful against multi-layer GNNs, suggesting that the linear dependency of privacy cost on network depth might be an overestimation. This observation aligns with the phenomenon known as “over-smoothing” [aaai/ChenLLLZS20] in GNNs, where node representations become increasingly homogeneous as network depth increases and consequently making membership inference more challenging. This homogenization effect might actually provide inherent privacy benefits due to the contractive nature of GNN aggregation operations. This observation motivates our central research question:
@@
In this work, we answer this question affirmatively. Prior perturbed message-passing mechanisms [sp/0011L0022, pakdd/IftikharW21, sigmod/DayLL16] assume that privacy loss grows linearly with depth, yet empirical results show that deeper GNNs can be less vulnerable to membership inference. We attribute this phenomenon to the contractive nature of common aggregation operators. In theory, we analyze this contractive property underlying over-smoothing, which leads to bounded sensitivity, so the privacy budget converges with the number of layers $K$ (see Figure 1) instead of growing linearly. This motivates $\mathsf{CARIBOU}$, a privacy-preserving GNN framework that enforces contractiveness to mirror real-world GNN behaviors while achieving convergent privacy budget through privacy amplification.
@@ I-A Overview of Convergent Privacy Analysis
For a perturbed message-passing GNN with $K$ layers, the standard approach analyzes the privacy loss at each step and then applies the DP composition theorem to aggregate the total privacy cost. This approach is common in existing privacy analyses of perturbed message-passing GNNs [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23, nips/chien2023differentially], resulting in a privacy budget that scales linearly with $K$, specifically $\epsilon=O(K/\sigma^{2})+O(\sqrt{K/\sigma})$. Thus, as $K$ increases, the amount of injected noise $\sigma$ must also grow, leading to degraded utility particularly when a small $\epsilon$ is desired; see Section III for a motivating empirical study.
Inspired by recent advances in privacy amplification [nips/ChourasiaYS21, icml/BokShifted24] through hidden states and contractive iterative processes, we observe that a similar amplification effect can be exploited in GNNs. Here, contractiveness refers to the property that the distance between two inputs is reduced after applying the operation, implying reduced distinguishability of outputs from a privacy perspective. The potential privacy amplification in GNNs arises from the following two observations: (1) GNNs typically do not expose intermediate node embeddings during training or inference, focusing only on the final node representations. (2) Standard message-passing operations, such as those used in Graph Convolutional Networks (GCNs) [iclr/KipfW17] (the dominant model in practice and in empirical studies [sp/Wulinkteller22]), are inherently contractive, a property that also underlies the over-smoothing phenomenon [aaai/ChenLLLZS20]. We theoretically validate this insight by showing that the privacy loss of a $K$-layer perturbed message-passing process with contractive layers satisfies a convergent privacy budget. Specifically, instead of growing linearly with $K$, we show that the privacy budget follows a convergent form: $\epsilon=O\left(\frac{(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})}{(1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}})}\right),$ where $C_{\mathsf{L}}$ is the Lipschitz constant of the message-passing operator (see Theorem 3 for details). Our improved privacy analysis is achieved by recasting the multi-layer perturbed GNN as a Contractive Noisy Iteration (CNI) process [icml/BokShifted24] and applying the privacy convergence results established for CNIs.
To leverage this analysis in practice, we design a simple yet effective Contractive Graph Layer (CGL) that enforces the contractiveness required for our theoretical guarantees while maintaining model expressivity. The CGL layer builds upon standard GCN-style aggregation, augmented with residual connections [icml/ChenWHDL20] and mean aggregation normalization [alon2020bottleneck], ensuring expressiveness across many layers without exposing sensitive edge information.
@@ I-B $\mathsf{CARIBOU}$: Framework and Evaluation
To evaluate $\mathsf{CARIBOU}$, we conduct extensive experiments over nine graph datasets, including commonly-used real-world datasets and synthetic chain-structured datasets for developing configurable interaction ranges. The experimental results demonstrate that $\mathsf{CARIBOU}$ improves non-trivial utility over standard graph and chain-structured datasets. Compared with several SOTA baselines, $\mathsf{CARIBOU}$’s EDP and NDP show significant utility improvements, especially in high privacy regimes, and reasonable computational overhead. Table I presents a comprehensive comparison, which is explained in Section VIII. Ablation studies are provided to understand the relation between privacy-utility hops and various ranges of graph, and the choice of hyper-parameters of CGL. In addition to privacy verification, we perform auditing experiments based on two membership inference attacks [sp/Wulinkteller22, tpsisa/OlatunjiNK21], demonstrating $\mathsf{CARIBOU}$’s robustness.
@@
Framework Mechanism Complexity per Layer Calibrated Noise ($\sigma$) EDP Utility NDP Utility PertGraph [ccs/KolluriBHS22, sp/0011L0022] Graph perturbation $O(|{\bm{V}}|^{2})$ $\propto 1$ DPDGC [nips/chien2023differentially] Decoupled graph with perturbation $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ GAP [uss/sajadmanesh2023gap] Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ $\mathsf{CARIBOU}$ Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{\min(K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}})}$
@@ II-A Message Passing Graph Neural Networks
Graph Neural Networks (GNNs) are a class of neural networks that operate on graph-structured data. Most GNNs follow the message-passing paradigm [feng2022powerful], where nodes iteratively aggregate information from their neighbors to update their representations.
@@ II-A1 Message Passing Layers
where $\sigma$ is a non-linear activation function, ${\mathcal{N}}(u)$ is the set of neighbors of node $u$, $\phi$ is a function that computes the message from node $v$ to node $u$, $\oplus$ represents the aggregation function that processes all messages from the neighbors of node $u$, and $\psi$ is a function that updates the node feature vector of node $u$ with the aggregated messages. Examples of message passing GNNs, such as GCN [iclr/kipf2017semi] and its variant, are in Appendix -A.
@@ II-A2 Applications of Message Passing GNNs
Message passing GNNs leverage GNN layers to iteratively refine node representations, which are then employed in tasks like node classification [ccs/KolluriBHS22], link prediction [www/RaoKR24], and graph classification [icdm/LiW024]. Multi-layer GNNs like deep GNNs [iclr/ThurlemannR23, nips/WuJWMGS21] are especially suitable to process long-range graphs [nips/DwivediRGPWLB22, corr/abs-2406-03386, tmlr/TonshoffRRG24] by capturing dependencies between distant nodes, which is crucial for tasks like molecular property prediction [bioinformatics/MaBRHXXYH22], protein interaction modeling [bmcbi/ZhongHXLQY22], and complex node interaction modeling [feng2022powerful].
@@ II-B Differential Privacy for GNNs
Definition 1 (Differential Privacy [tcc/DworkMNS06]).
@@
Intuitively, DP [tcc/DworkMNS06] theoretically quantifies the privacy of a model by measuring the indistinguishability of the outputs of a mechanism $\mathcal{M}$ on two adjacent datasets $D$ and $D^{\prime}$. It can be classified into bounded DP and unbounded DP depending on the construction of $D^{\prime}$, where the former is by replacing a data instance of $D$ and the latter is by addition / removal of a data sample of $D$. The privacy budget $\epsilon$ is smaller representing a stronger privacy guarantee, while $\delta$ is a slackness quantity that relaxes the pure DP constraint.
@@ II-B1 Privacy Definition on Graphs
Definition 2 (Edge-level adjacency [pvldb/KarwaRSY11]).
@@
Definition 3 (Node-level adjacency [pvldb/KarwaRSY11]).
@@ II-B2 Perturbed Message Passing with DP
To incorporate differential privacy into GNNs, one can add noise to the message passing layer, following the perturbed message passing approach [uss/sajadmanesh2023gap]. Given a graph $\bm{G}$} and message passing function $\mathrm{MP}$, we define a sequence $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ of node feature matrices by:
@@ II-B3 Privacy Accounting
Privacy accounting process analyzes the total privacy budget for the composition of several (adaptive) private algorithms. A common approach for analyzing the Gaussian mechanism in perturbed message passing is through Rényi differential privacy (RDP) [csf/mironov2017renyi] and its composition theorem.
Definition 5 (Rényi differential privacy [csf/mironov2017renyi]).
@@
Theorem 1 (Composition of RDP [csf/mironov2017renyi]).
@@
In this work, we present our privacy results in terms of RDP for ease of interpretation, while our underlying analysis employs the tighter $f$-DP framework. This analysis leverages recent advances in privacy amplification techniques [nips/SchuchardtSKG24, icml/BokShifted24] to achieve stronger privacy guarantees. The technical details of our convergent privacy analysis are discussed in §III-B.
@@ III Private Multi-layer GNNs Initiative
Multi-layer GNNs [iclr/ThurlemannR23, nips/WuJWMGS21] are vital in tasks like modeling molecular structures, where some properties depend on long range interactions [nips/DwivediRGPWLB22, corr/abs-2406-03386, tmlr/TonshoffRRG24] of the nodes [zhang2022deep, li2023neural]. Specifically, they require messages flowing across multiple hops before reaching a target node through stacking multiple layers to exchange information across $K$-hop neighborhoods. However, ensuring privacy in multi-layer GNNs poses key challenges. In this section, we outline and illustrate these challenges via a motivating case study (Section III-A), ultimately motivating our new design insights (Section III-B) based on contractive message passing.
@@ III-A1 A Case Study on GAP’s Performance on Learning Long-Range Interactions
To further investigate this phenomenon, we evaluate $K$-layer GAP on the chain-structured dataset (see Figure 2), which is adopted in [nips/GuC0SG20, liu2021eignn] to examine long-range interaction learning capabilities. Specifically, this dataset creates a controlled environment for evaluating GNN performance on long-range interactions. A model without learning from graph structure such as MLP would fail since most nodes in the dataset have zero-valued feature vectors regardless of their chain type. For a message-passing GNN to correctly classify nodes positioned $K$ hops away from the informative first node, it must perform at least $K$ propagation steps to transfer the meaningful features across the chain as shown in Figure 2. This requirement becomes particularly challenging in the private setting as noise must be injected after each message passing layer, potentially overwhelming the signal being propagated.

### v3

@@ I Introduction
Graph neural networks (GNNs) [36, 25, 72], designed for operating over structural data, have achieved success in various domains, including social networks [55, 96] and recommendation systems [10, 80, 83]. At their core, many GNN architectures are built upon the message-passing paradigm, where node representations are iteratively updated by aggregating information from their neighbors. However, graph structures often encode sensitive information about relationships and attributes. As a result, GNNs are vulnerable to privacy attacks, including membership inference [78, 90, 29] and attribute inference [99, 66]. These vulnerabilities highlight the urgent need for robust privacy protection mechanisms in graph learning.
@@
Differential privacy (DP) [49, 76, 17] has emerged as a foundational framework to provide formal guarantees against data leakage over the past two decades, with widespread applications in machine learning [68, 92, 88], synthetic data generation [65, 41, 47, 54], and beyond. In the context of GNNs, recent works [57, 82] have advanced privacy protection through edge-level DP (EDP) [77] and node-level DP (NDP) [30, 16] guarantees. The primary approach employs perturbed message passing, which injects calibrated Gaussian or Laplace noise into aggregation layers to protect the edge or node memberships in the training graph. While these approaches have provided formal privacy guarantees, they share a critical limitation: the privacy loss grows linearly with the number of layers $K$ or graph hops. In other words, each additional aggregation layer compounds the privacy cost, ultimately requiring large amounts of noise to maintain a reasonable level of privacy protection. This, in turn, severely degrades model utility.
@@
Recent advances have shown that multi-layer GNNs, especially deeper GCNs, are essential in capturing complex relationships [22] and analyzing graphs with long-range interactions [18, 9, 84, 70, 69]. In fact, informative long-range interactions exist in a lot of real-world datasets. For instance, in large biological networks, long-range dependencies influence protein functions, requiring more than 10 hops of message passing [39]. In social networks, privacy-sensitive relationships propagate through multi-hop neighborhoods [50, 43]. As reported in [39], increasing the network depth leads to a substantial improvement in accuracy from $72.5\%$ to $88.2\%$. However, the aforementioned linear dependence on $K$ is particularly challenging for multi-layer GNNs as larger $K$ leads to larger privacy parameter $\epsilon$, a.k.a weak privacy guarantee.
@@
Interestingly, empirical studies [78] have shown that membership inference attacks are not particularly more successful against multi-layer GNNs, suggesting that the linear dependency of privacy cost on network depth might be an overestimation. This observation aligns with the phenomenon known as “over-smoothing” [8] in GNNs, where node representations become increasingly homogeneous as network depth increases and consequently making membership inference more challenging. This homogenization effect might actually provide inherent privacy benefits due to the contractive nature of GNN aggregation operations. This observation motivates our central research question:
@@
In this work, we answer this question affirmatively. Prior perturbed message-passing mechanisms [77, 30, 16] assume that privacy loss grows linearly with depth, yet empirical results show that deeper GNNs can be less vulnerable to membership inference. We attribute this phenomenon to the contractive nature of common aggregation operators. In theory, we analyze this contractive property underlying over-smoothing, which leads to bounded sensitivity, so the privacy budget converges with the number of layers $K$ (see Figure 1) instead of growing linearly. This motivates $\mathsf{CARIBOU}$, a privacy-preserving GNN framework that enforces contractiveness to mirror real-world GNN behaviors while achieving convergent privacy budget through privacy amplification.
@@ I-A Overview of Convergent Privacy Analysis
For a perturbed message-passing GNN with $K$ layers, the standard approach analyzes the privacy loss at each step and then applies the DP composition theorem to aggregate the total privacy cost. This approach is common in existing privacy analyses of perturbed message-passing GNNs [57, 82, 13], resulting in a privacy budget that scales linearly with $K$, specifically $\epsilon=O(K/\sigma^{2})+O(\sqrt{K/\sigma})$. Thus, as $K$ increases, the amount of injected noise $\sigma$ must also grow, leading to degraded utility particularly when a small $\epsilon$ is desired; see Section III for a motivating empirical study.
Inspired by recent advances in privacy amplification [14, 6] through hidden states and contractive iterative processes, we observe that a similar amplification effect can be exploited in GNNs. Here, contractiveness refers to the property that the distance between two inputs is reduced after applying the operation, implying reduced distinguishability of outputs from a privacy perspective. The potential privacy amplification in GNNs arises from the following two observations: (1) GNNs typically do not expose intermediate node embeddings during training or inference, focusing only on the final node representations. (2) Standard message-passing operations, such as those used in Graph Convolutional Networks (GCNs) [36] (the dominant model in practice and in empirical studies [78]), are inherently contractive, a property that also underlies the over-smoothing phenomenon [8]. We theoretically validate this insight by showing that the privacy loss of a $K$-layer perturbed message-passing process with contractive layers satisfies a convergent privacy budget. Specifically, instead of growing linearly with $K$, we show that the privacy budget follows a convergent form: $\epsilon=O\left(\frac{(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})}{(1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}})}\right),$ where $C_{\mathsf{L}}$ is the Lipschitz constant of the message-passing operator (see Theorem 3 for details). Our improved privacy analysis is achieved by recasting the multi-layer perturbed GNN as a Contractive Noisy Iteration (CNI) process [6] and applying the privacy convergence results established for CNIs.
To leverage this analysis in practice, we design a simple yet effective Contractive Graph Layer (CGL) that enforces the contractiveness required for our theoretical guarantees while maintaining model expressivity. The CGL layer builds upon standard GCN-style aggregation, augmented with residual connections [11] and mean aggregation normalization [2], ensuring expressiveness across many layers without exposing sensitive edge information.
@@ I-B $\mathsf{CARIBOU}$: Framework and Evaluation
To evaluate $\mathsf{CARIBOU}$, we conduct extensive experiments over nine graph datasets, including commonly-used real-world datasets and synthetic chain-structured datasets for developing configurable interaction ranges. The experimental results demonstrate that $\mathsf{CARIBOU}$ improves non-trivial utility over standard graph and chain-structured datasets. Compared with several SOTA baselines, $\mathsf{CARIBOU}$’s EDP and NDP show significant utility improvements, especially in high privacy regimes, and reasonable computational overhead. Table I presents a comprehensive comparison, which is explained in Section VIII. Ablation studies are provided to understand the relation between privacy-utility hops and various ranges of graph, and the choice of hyper-parameters of CGL. In addition to privacy verification, we perform auditing experiments based on two membership inference attacks [78, 53], demonstrating $\mathsf{CARIBOU}$’s robustness.
@@
Framework Mechanism Complexity per Layer Calibrated Noise ($\sigma$) EDP Utility NDP Utility PertGraph [38, 77] Graph perturbation $O(|{\bm{V}}|^{2})$ $\propto 1$ DPDGC [13] Decoupled graph with perturbation $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ GAP [59] Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ $\mathsf{CARIBOU}$ Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{\min(K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}})}$
@@ II-A Message Passing Graph Neural Networks
Graph Neural Networks (GNNs) are a class of neural networks that operate on graph-structured data. Most GNNs follow the message-passing paradigm [22], where nodes iteratively aggregate information from their neighbors to update their representations.
@@ II-A1 Message Passing Layers
where $\sigma$ is a non-linear activation function, ${\mathcal{N}}(u)$ is the set of neighbors of node $u$, $\phi$ is a function that computes the message from node $v$ to node $u$, $\oplus$ represents the aggregation function that processes all messages from the neighbors of node $u$, and $\psi$ is a function that updates the node feature vector of node $u$ with the aggregated messages. Examples of message passing GNNs, such as GCN [35] and its variant, are in Appendix -A.
@@ II-A2 Applications of Message Passing GNNs
Message passing GNNs leverage GNN layers to iteratively refine node representations, which are then employed in tasks like node classification [38], link prediction [56], and graph classification [40]. Multi-layer GNNs like deep GNNs [69, 84] are especially suitable to process long-range graphs [18, 9, 70] by capturing dependencies between distant nodes, which is crucial for tasks like molecular property prediction [46], protein interaction modeling [100], and complex node interaction modeling [22].
@@ II-B Differential Privacy for GNNs
Definition 1 (Differential Privacy [19]).
@@
Intuitively, DP [19] theoretically quantifies the privacy of a model by measuring the indistinguishability of the outputs of a mechanism $\mathcal{M}$ on two adjacent datasets $D$ and $D^{\prime}$. It can be classified into bounded DP and unbounded DP depending on the construction of $D^{\prime}$, where the former is by replacing a data instance of $D$ and the latter is by addition / removal of a data sample of $D$. The privacy budget $\epsilon$ is smaller representing a stronger privacy guarantee, while $\delta$ is a slackness quantity that relaxes the pure DP constraint.
@@ II-B1 Privacy Definition on Graphs
Definition 2 (Edge-level adjacency [33]).
@@
Definition 3 (Node-level adjacency [33]).
@@ II-B2 Perturbed Message Passing with DP
To incorporate differential privacy into GNNs, one can add noise to the message passing layer, following the perturbed message passing approach [59]. Given a graph $\bm{G}$} and message passing function $\mathrm{MP}$, we define a sequence $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ of node feature matrices by:
@@ II-B3 Privacy Accounting
Privacy accounting process analyzes the total privacy budget for the composition of several (adaptive) private algorithms. A common approach for analyzing the Gaussian mechanism in perturbed message passing is through Rényi differential privacy (RDP) [49] and its composition theorem.
Definition 5 (Rényi differential privacy [49]).
@@
Theorem 1 (Composition of RDP [49]).
@@
In this work, we present our privacy results in terms of RDP for ease of interpretation, while our underlying analysis employs the tighter $f$-DP framework. This analysis leverages recent advances in privacy amplification techniques [60, 6] to achieve stronger privacy guarantees. The technical details of our convergent privacy analysis are discussed in §III-B.
@@ III Private Multi-layer GNNs Initiative
Multi-layer GNNs [69, 84] are vital in tasks like modeling molecular structures, where some properties depend on long range interactions [18, 9, 70] of the nodes [95, 42]. Specifically, they require messages flowing across multiple hops before reaching a target node through stacking multiple layers to exchange information across $K$-hop neighborhoods. However, ensuring privacy in multi-layer GNNs poses key challenges. In this section, we outline and illustrate these challenges via a motivating case study (Section III-A), ultimately motivating our new design insights (Section III-B) based on contractive message passing.
@@ III-A1 A Case Study on GAP’s Performance on Learning Long-Range Interactions
To further investigate this phenomenon, we evaluate $K$-layer GAP on the chain-structured dataset (see Figure 2), which is adopted in [25, 44] to examine long-range interaction learning capabilities. Specifically, this dataset creates a controlled environment for evaluating GNN performance on long-range interactions. A model without learning from graph structure such as MLP would fail since most nodes in the dataset have zero-valued feature vectors regardless of their chain type. For a message-passing GNN to correctly classify nodes positioned $K$ hops away from the informative first node, it must perform at least $K$ propagation steps to transfer the meaningful features across the chain as shown in Figure 2. This requirement becomes particularly challenging in the private setting as noise must be injected after each message passing layer, potentially overwhelming the signal being propagated.